# HoloDx: Knowledge- and Data-Driven Multimodal Diagnosis of Alzheimer’s Disease

### v1

@@ 4.2.4 Factor Analysis
Shapley analysis [38] is performed on the ADNI, AIBL, and RENJI test sets to identify the numerical features that most significantly influenced the model’s diagnostic decisions (Fig. 3). Across all three datasets, the MMSE score consistently ranks among the top two most influential features. In the AIBL dataset, the Immediate Recall score (LIMMTOTAL), a subcomponent of MMSE, is ranked as the most important feature, and it uniquely appears in this dataset. For both ADNI and AIBL, MMSE, Thyroid Stimulating Hormone, Vitamin B12 levels, and the presence of APOE4 alleles are selected consistently among the top ten factors. These findings align with clinical studies that emphasize the strong association of MMSE scores and other key biomarkers with cognitive impairment and AD diagnosis.
In the RENJI dataset, Auditory Verbal Learning Test (AVLT) scores and its subscore emerge in the top ten factors, closely aligned with the knowledge provided by physicians and clinical studies [39]. AVLT is a test designed to assess auditory memory, which is valuable for detecting early-stage cognitive decline. Another interesting finding is the appearance of BMI and weight in the top ten, consistent with other clinical studies linking these factors to cognitive impairment [40, 41].
@@ 4.2.5 Biomarker-Level Validation
To improve clinical accessibility, we use easily available factors that do not involve pathological AD biomarkers, and conduct biomarker-level verification to validate model effectiveness. As shown in the raincloud plots in Fig. 4, our predictions correlate significantly with PET biomarkers (AV45, FDG, PIB) and CSF biomarkers (PTAU, TAU, Abeta) in the ADNI cohort, as well as with PET AV45 and Serum PTAU217 biomarkers in the RENJI cohort, demonstrating statistically significant differences between NC and CI groups (p-values 0.001). Notably, PTAU217 exhibits greater diagnostic efficacy than Serum PTAU181 and Abeta42, further supporting its role in detecting cerebral amyloidosis and predicting cognitive decline [42].

### v2

@@ 4.2.4 Factor Analysis
Shapley analysis [46] is performed on the ADNI, AIBL, and RENJI test sets to identify the numerical features that most significantly influenced the model’s diagnostic decisions (Fig. 4). Across all three datasets, the MMSE score consistently ranks among the most influential features. For both ADNI and AIBL, MMSE, Thyroid Stimulating Hormone, Vitamin B12 levels, and the presence of APOE4 alleles are selected consistently among the top ten factors. These findings align with clinical studies that emphasize the strong association of MMSE scores and other key biomarkers with cognitive impairment and AD diagnosis.
In the RENJI dataset, Auditory Verbal Learning Test (AVLT) scores and their subscores emerge in the top ten factors, closely aligned with the knowledge provided by physicians and clinical studies [47]. AVLT is a test designed to assess auditory memory, which is valuable for detecting early-stage cognitive decline. Notably, elevated serum aspartate aminotransferase (AST) levels, previously associated with AD in comparative studies [48], and Body Mass Index (BMI), linked to cognitive impairment in other research [49, 50], also featured prominently in the RENJI top ten features.
Factor | LLM | Knowledge |
MMSE | GPT4o [51] | A lower MMSE score indicates poorer cognitive function, which is often seen in individuals with Alzheimer’s disease. |
Llama3 [52] | Lower MMSE scores are associated with poorer cognitive function and an increased risk of AD. Low MMSE scores may indicate a higher risk of AD. | |
Med-llama3 [53] | The Mini-Mental State Examination (MMSE) is a brief cognitive test that is widely used in clinical practice. Its diagnostic accuracy for Alzheimer’s disease (AD) in the memory clinic population is unclear. We aimed to determine the diagnostic accuracy of the MMSE in a memory clinic population. The study was a secondary analysis of the Australian Imaging, Biomarkers and Lifestyle (AIBL) Study of Ageing. | |
HuatuoGPT-V [13] | A lower MMSE score indicates poorer cognitive performance, which may be associated with an increased likelihood of AD diagnosis. However, the relationship between MMSE scores and AD diagnosis can also depend on other factors such as age, education level, and severity of symptoms. | |
APOE | GPT4o [51] | The presence of the APOE-4 allele increases an individual’s genetic risk for developing Alzheimer’s disease (AD), influencing both the likelihood and the potential age of onset. |
Llama3 [52] | Carrying the APOE-4 allele is a significant genetic risk factor for Alzheimer’s disease (AD); individuals with one or two copies of the APOE-e4 gene have an increased likelihood of developing AD. | |
Med-llama3 [53] | The apolipoprotein E (APOE) gene is a major risk factor for Alzheimer’s disease (AD). The APOE 4 allele is associated with a 3- to 4-fold increased risk of developing AD, while the APOE 2 allele is associated with a reduced risk. The aim of this study was to systematically review the evidence on the relationship between APOE and AD diagnosis, and to perform a meta-analysis of the diagnostic accuracy of APOE genotyping in AD. | |
HuatuoGPT-V [13] | The APOE gene has three main variants: 2, 3 (most common), and 4. The presence of the 4 allele is associated with an increased risk for Alzheimer’s disease (AD) diagnosis; individuals homozygous or heterozygous for the 4 variant have a higher likelihood of developing AD compared to those without this genetic factor. |
@@ 4.2.5 Biomarker-Level Validation
To enhance clinical accessibility, our model relies on readily available non-pathological factors and validates its effectiveness through biomarker-level analysis. As shown in the raincloud plots (Fig. 5), we visualize the distributions of biomarker values for individuals predicted as cognitively normal (NC, dark blue) and cognitively impaired (CI, light blue). Each raincloud plot comprises three elements: (i) the “cloud”, representing a smoothed density curve of the value distribution; (ii) the “rain”, showing individual data points jittered horizontally to reflect raw cohort-level observations; (iii) the boxplot, summarizing the median, quartiles, and outliers. This integrated visualization reveals strong concordance between model predictions and known neuropathological markers. In the ADNI dataset, predictions correlate with PET biomarkers (AV45, FDG, PIB) and CSF biomarkers (PTAU, TAU, Abeta), while in the RENJI cohort, they align with PET AV45 and serum pTau217 levels. Notably, both predicted classes and biomarker distributions show clear separation between NC and CI groups (all inter-group comparisons: ), confirming the biological plausibility of the model’s outputs. Among these, serum pTau217 exhibits superior discriminative power compared to pTau181 and Abeta42, consistent with its established utility in detecting cerebral amyloidosis and predicting cognitive decline [54].
4.2.6 Comparison of Different LLMs for Knowledge Generation
We conduct a comprehensive evaluation of several state-of-the-art LLMs, including GPT-4o [51], HuatuoGPT-Vision [13], Llama3 [52], and Med-Llama3 [53]. Qualitative examples of knowledge generation for representative factors (MMSE and APOE) are detailed in Table 11. HuatuoGPT-V consistently provides the most clinically relevant and comprehensive descriptions of both MMSE and APOE factors, effectively incorporating essential clinical context, individual variability, and explicit differentiation among genetic variants. GPT-4o, although succinct and accurate, lacks certain contextual nuances and variant-specific details essential for clinical applications. Meanwhile, Med-Llama3 and Llama3 exhibit notable limitations, including overly academic phrasing and insufficient detail, which significantly reduces their practical utility.
We further employ automated metrics provided by the third-party LLM Deepseek [55] to evaluate the quality of the generated knowledge, supplemented by expert spot-check validation. This assessment focuses on their capability to generate clinically accurate descriptions of disease-related factors across four critical dimensions: terminology accuracy, clinical relevance, practical utility, and consistency. Quantitative comparisons of the models’ knowledge element outputs are systematically summarized in Table 12. Although GPT-4o achieved the highest quantitative performance, its computational demands made it unsuitable for integration into our lightweight diagnostic framework. In contrast, HuatuoGPT-V demonstrated superior clinical relevance and consistency compared to Llama3 and Med-Llama3 (1569% improvement). Furthermore, additional experiments on the AIBL dataset (Table 13) confirmed minimal impact on diagnostic performance (% AUC difference and the same accuracy) between GPT-4o and HuatuoGPT-V, validating our selection of HuatuoGPT-V for its balanced performance and practical deployability.
These combined quantitative results and qualitative assessments underscore the superiority of HuatuoGPT-V, substantiating its optimal balance of clinical precision, explanatory depth, and deployment efficiency.
Model | Evaluation Dimensions | Total | |||
Terminology | Relevance | Utility | Consistency | ||
GPT-4o [56] | 98.6 | 99.8 | 96.2 | 99.1 | 98.5 |
Llama3 [52] | 83.2 | 71.4 | 63.8 | 78.3 | 73.8 |
Med-Llama3 [53] | 27.5 | 18.3 | 12.6 | 23.7 | 20.0 |
HuatuoGPT-V [13] | 91.7 | 93.5 | 84.9 | 87.4 | 89.7 |
1Weighting: Terminology (35%), Relevance (30%), Utility (25%), Consistency (10%) |
2Total score = (Dimension score Dimension weight) |
Task | ACC | AUC | SEN | SPE |
Llama3 [52] | 86.87 | 81.11 | 94.37 | 67.86 |
HuatuoGPT-V [13] | 91.92 | 88.23 | 97.14 | 79.31 |
GPT-4o [56] | 91.92 | 89.24 | 95.71 | 82.76 |
Task | ACC | AUC | SEN | SPE |
LLM (text-only) | 75.76 | 60.64 | 97.14 | 24.14 |
VLLM (text + 2D slice) | 76.77 | 60.01 | 98.60 | 21.43 |
HoloDx (ours) | 91.92 | 88.23 | 97.14 | 79.31 |
Method | Training Time | GPU Mem | Parameters | Inference Time |
/Epoch (min) | (GB) | (MB) | /Sample (ms) | |
AD-Trans [9] | 2.6 | 4.7 | 201.5 | 135.2 |
Alifuse [7] | 7.4 | 14.1 | 608.7 | 178.1 |
Backbone (our BB) | 6.4 | 9.7 | 228.2 | 223.6 |
BB+KL | 8.7 (+35.9%) | 12.1 (+24.7%) | 276.0 (+20.9%) | 231.7 (+3.6%) |
BB+KL+MEM | 10.5 (+64.1%) | 13.1 (+35.0%) | 303.8 (+33.1%) | 232.9 (+4.2%) |
4.2.7 Direct LLM/VLLM Diagnosis vs. Specialized Multimodal Modeling
Our HoloDx is designed to address key limitations of generic LLM or vision-augmented (VLLM) methods by deeply integrating clinical expertise and imaging information. Specifically, the framework incorporates 3D brain images to capture structural and metabolic abnormalities, embeds domain-specific clinical data and knowledge aligned with AD pathology, and leverages LLMs as adaptive knowledge retrievers, rather than standalone diagnostic agents. This design enables our model to deliver clinically grounded, individualized interpretations while ensuring efficiency and generalization.
To validate the superiority of our specialized method, we conduct a comparative evaluation reported in Table 14. Pure text-based LLM diagnosis, which relies solely on symptom descriptions without imaging data, exhibits 16.16% lower accuracy compared to HoloDx. The VLLM achieves moderate performance (76.77% accuracy on AIBL test dataset) but remains constrained by its dependence on 2D visual features and general medical knowledge, lacking the capacity to model 3D neuroimaging patterns or domain-specific clinical context. In contrast, our multimodal design achieves 91.92% diagnostic accuracy, demonstrating a statistically significant improvement over generic LLM/VLLM methods.
4.2.8 Comparison of Model Efficiency and Complexity
To address concerns regarding computational and storage demands, we conducted comprehensive efficiency measurements across all methods using a standardized hardware platform under identical software conditions. As systematically quantified in Table 15, our complete framework incorporating both knowledge and memory injection modules exhibits a measured increase in resource requirements while maintaining clinically viable performance characteristics. The knowledge injection component introduces moderate overhead—35.9% longer training times and 24.7% higher memory consumption—primarily attributable to its cross-modal attention mechanisms and the computational cost of knowledge alignment during optimization. When augmenting this foundation with our memory injection module, we observe further increases in training time (64.1% total increase over our backbone) and parameter count (33.1% expansion), though crucially with only marginal impact on inference latency (4.2% increase). While the training time per epoch increases to 10.5 minutes and peak memory utilization reaches 13.1GB, these remain within practical boundaries for hospital computing infrastructure. The measured resource increases are clinically justified by our method’s 5.05% AUC improvement over baseline (Table 7(b). When compared to contemporary approaches, our solution demonstrates superior memory efficiency relative to Alifuse [7] while maintaining competitive training durations.